If velocity=4, the quantity of money= 20,000, and the price level =2.5 then the real value of output is....?

A) 2,000

B)200,000

C)12,500

D)32,000
Answer #1

As per quantity theory

M*V=P*Q

20000*4=2.5*Q

Q=80000/2.5

Q=32000 is real value of output